Operation Summary for September 16, 2009 - September 18, 2009 |
Wednesday September 16 The day shift began with Operations monitoring the Tevatron (TeV) store 7183 (22.27E30) and the stash (127.2E10), with the Antiproton Source (Pbar) stacking (6.7E10), and with startup operations and tuning in progress for Linac, Booster, the Main Injector (MI), Recycler, and with the NuMI and MiniBooNE experiments taking beam. During the day shift, NuMI personnel accessed their absorber hall to investigate dehumidifier issues; beam was re-established by noon. Operations terminated store 7183 at 1:45 PM. Proton only studies followed. The 6 mA antiproton stack was transferred to the Recycler and stacking began in earnest by 2:30 PM. Recycler, Pelletron, and MI personnel conducted beam studies. During the evening shift, the MI collimators were commissioned, which allowed the delivery of slip-stacked beam to Pbar and NuMI. TeV personnel conducted beam studies until 7:15 PM. The TeV separators were placed in conditioning mode and the TeV was left at 150 GeV. Recycler studies concluded at 8:30 PM. The Recycler stashed antiprotons. Booster suffered from beam position problems at 9:35 PM due to a high voltage cable problem. Operations called in Booster personnel to re-terminate the cable. The NuMI extraction kicker failed at 10:24 PM and would not reset. Operations called in EE Support personnel to investigate. Shift ended with experts addressing the problems with the Booster notcher and NuMI kicker. Thursday September 17 The midnight shift began with Operations monitoring the stash (29.3E10) and stack (26.8E10), but with no beam due to Booster notcher repairs. During the midnight shift, the notcher and NuMI extraction kicker repairs were completed by 1:30 and 3:30 AM respectively. The NuMI extraction kicker was pulsed for ~1.5 hours before re-establishing beam at 5:30 AM. The Recycler stashed when appropriate. During the day shift, NuMI kicker fluorinert water skid repairs were complete by 11:15 AM and NuMI beam was re-established. TeV personnel conducting beam studies until 11:30 AM. HEP shot setup followed; controls personnel investigated continuing BPM problems. Operations established store 7188 at 2:15 PM with an average initial luminosity of 134.13E30. During the evening shift, the Recycler stashed when appropriate. Operations monitored the vacuum in the MI-40 area and increased beam intensities as the vacuum improved. Friday September 18 The midnight shift began with Operations monitoring store 7188 (46.67E30) and the stash (154.6E10), with Pbar stacking (23.7E10), and with MiniBooNE and NuMI taking beam. During the midnight shift, the Recycler stashed when appropriate. Beam intensities were increased as the vacuum in MI-40 improved. The Plans for Friday and the Weekend A new store will go in this morning. MI personnel will continue to tune and increase beam intensities. The plans for today and the rest of the weekend are to stack, stash, and store. MiniBooNE beam will be turned off for vacuum work on Monday morning and will be off for about a day. |
|
More Information Comments and Suggestions |
| last modified 09/21/2009 email Fermilab |
| Security, Privacy, Legal | |